OpenStudy (anonymous):

6x2 – 15x = 0 a) x= 2 or x = 5.5 b) x= 0 or x= 2 c) x= 2 or x= 5 d) x= 0 or x= 2.5

OpenStudy (anonymous):

c

Parth (parthkohli):

3x(2x - 5) = 0 2x - 5 = 0/3x [if x is not 0] 2x - 5 = 0 2x = 5 x = 5/2 x can be both 0 or 2.5 Answer is d

Parth (parthkohli):

@Rohangrr please check, am I wrong or are you?